# EvoAgentX: An Automated Framework for Evolving Agentic Workflows

### v1

@@ 1 Introduction
Multi-agent systems (MAS) are emerging as a powerful paradigm for orchestrating large language models (LLMs) and specialized tools to solve complex tasks collaboratively Hong et al. (2023); Gao et al. (2024). By coordinating multiple agents with distinct capabilities, such as retrieval, planning, reasoning, or code generation, MAS decompose intricate problems into controllable subtasks and assign them to agents capable of solving them Yuan et al. (2024); Zhang et al. (2025a). This flexible and modular architecture makes MAS well-suited for addressing complex real-world problems. As a result, MAS have been widely deployed in applications such as multi-hop question answering Hong et al. (2023), software engineering automation Li et al. (2023), code generation Liu et al. (2025), mathematical problem solving Gao et al. (2024), and dialogue systems Shi et al. (2024).
@@
In this paper, we propose EvoAgentX, an open-source platform designed for the automated generation and evolutionary optimization of multi-agent workflows. As illustrated in Figure 1, EvoAgentX employs a modular architecture consisting of five core layers: basic components, agent, workflow, evolving, and evaluation. The central feature of EvoAgentX is the evolving layer, which seamlessly integrates three state-of-the-art optimization algorithms, TextGrad Yuksekgonul et al. (2024), AFlow Zhang et al. (2024b), and MIPRO Opsahl-Ong et al. (2024), to iteratively refine agent prompts, tool configurations, and workflow topologies. We evaluate EvoAgentX comprehensively across diverse benchmarks, including HotPotQA for multi-hop reasoning Yang et al. (2018), MBPP for code generation Austin et al. (2021a), MATH for mathematical problem-solving Hendrycks et al. (2021), and the GAIA benchmark for real-world multi-agent tasks Mialon et al. (2023). Experimental results demonstrate that EvoAgentX achieves substantial performance improvements through dynamic workflow evolution, including improvements of 7.44% in HotPotQA F1, 10.00% in MBPP pass@1, 10.00% in MATH solve accuracy, and up to 20.00% overall accuracy on GAIA.
The contributions of our paper are as follows:
@@ 3.3 Workflow Layer
The workflow layer is a core component of EvoAgentX, supporting the construction, orchestration, and execution of multi-agent workflows in a structured and flexible manner. It provides a formal representation for capturing task dependencies, execution flows, and communication between agents. Each workflow is modeled as a directed graph:
@@ 3.4 Evolving Layer
(1) The agent optimizer aims to refine agent prompt templates, tool configurations, and action strategies to enhance individual agent performance across diverse tasks. Formally, for an agent parameterized by its prompt and configuration , the optimizer seeks to compute
where denotes the agent-level optimization operator that updates prompts and configurations based on evaluation feedback, and denotes evaluation feedback. The TextGrad Yuksekgonul et al. (2024) and MIPRO Opsahl-Ong et al. (2024) optimizers are employed for agent optimization, jointly applying gradient-based prompt tuning, in-context learning, and preference-guided refinement to iteratively align prompts, tool configurations, and agent outputs with task-specific objectives based on performance signals.
@@
where denotes the workflow-level optimization operator that updates the graph structure based on evaluation feedback, and denotes evaluation feedback. The SEW Liu et al. (2025) and AFlow Zhang et al. (2024b) optimizers are employed for workflow optimization, iteratively restructuring workflow graphs by reordering nodes, modifying dependencies, and exploring alternative execution strategies guided by task performance signals and convergence criteria.
@@ 5 Conclusion
We present EvoAgentX, an open-source platform that automates the generation, execution, and evolutionary optimization of multi-agent workflows. It addresses key limitations of existing frameworks by eliminating the need for manual workflow design and providing support for dynamic, task-specific optimization. By integrating multiple optimization algorithms, including TextGrad, AFlow, and MIPRO, EvoAgentX enables the iterative refinement of agent prompts, tool configurations, and workflow topologies with minimal manual intervention. Experiments on diverse benchmarks, such as HotPotQA, MBPP, MATH and GAIA, demonstrate that EvoAgentX consistently achieves substantial performance improvements in multi-hop reasoning, code generation, mathematical problem solving, and real-world multi-agent applications. In the future, we will extend EvoAgentX with plug-and-play prompt optimization, richer tool integration, and long-term memory with retrieval-augmented generation (RAG) support to further enhance agent adaptability and contextual awareness. We also plan to explore advanced evolution strategies, including MASS, EvoPrompt, and Darwin, to advance the state of dynamic multi-agent optimization.

### v2

@@ 1 Introduction
Multi-agent systems (MAS) are emerging as a powerful paradigm for orchestrating large language models (LLMs) and specialized tools to solve complex tasks collaboratively Hong et al. (2023); Gao et al. (2024); Fang et al. (2025). By coordinating multiple agents with distinct capabilities, such as planning, reasoning, or code generation, MAS decompose intricate problems into controllable subtasks and assign them to agents capable of solving them Yuan et al. (2024); Zhang et al. (2025a). This flexible and modular architecture makes MAS well-suited for addressing complex real-world problems. As a result, MAS have been widely deployed in applications such as multi-hop question answering Hong et al. (2023), software engineering automation Li et al. (2023), code generation Liu et al. (2025), mathematical problem solving Gao et al. (2024), and dialogue systems Shi et al. (2024).
@@
To this end, we propose EvoAgentX, an open-source platform designed for the automated generation and evolutionary optimization of multi-agent workflows. As illustrated in Figure 1, EvoAgentX employs a modular architecture consisting of five core layers: basic components, agent, workflow, evolving, and evaluation. The central feature of EvoAgentX is the evolving layer, which seamlessly integrates three state-of-the-art optimization algorithms, TextGrad Yuksekgonul et al. (2024), AFlow Zhang et al. (2024b), and MIPRO Opsahl-Ong et al. (2024), to iteratively refine agent prompts, tool configurations, and workflow topologies. We evaluate EvoAgentX comprehensively across diverse benchmarks, including HotPotQA for multi-hop reasoning Yang et al. (2018), MBPP for code generation Austin et al. (2021a), MATH for mathematical problem-solving Hendrycks et al. (2021), and the GAIA benchmark for real-world multi-agent tasks Mialon et al. (2023). Experimental results demonstrate that EvoAgentX achieves substantial performance improvements through dynamic workflow evolution, including improvements of 7.44% in HotPotQA F1, 10.00% in MBPP pass@1, 10.00% in MATH solve accuracy, and up to 20.00% overall accuracy on GAIA.
The contributions of our demo are as follows:
@@ 3.3 Workflow Layer
The workflow layer is another core component of EvoAgentX, supporting the construction, orchestration, and execution of multi-agent workflows in a structured and flexible manner. It provides a formal representation for capturing task dependencies, execution flows, and communication between agents. Each workflow is modeled as a directed graph:
@@ 3.4 Evolving Layer
(1) The agent optimizer aims to refine agent prompt templates, tool configurations, and action strategies to enhance each agent’s performance across diverse tasks. Formally, for an agent parameterized by its prompt and configuration , the optimizer seeks to compute
where denotes the agent-level optimization operator that updates prompts and configurations based on evaluation feedback, and denotes evaluation feedback. The TextGrad Yuksekgonul et al. (2024) and MIPRO Opsahl-Ong et al. (2024) optimizers are employed for agent optimization, jointly applying gradient-based prompt tuning, in-context learning, and preference-guided refinement to iteratively align prompts, tool configurations, and agent outputs with task-specific objectives based on the performance signal.
@@
where denotes the workflow-level optimization operator that updates the graph structure based on evaluation feedback, and denotes evaluation feedback. The SEW Liu et al. (2025) and AFlow Zhang et al. (2024b) optimizers are employed for workflow optimization, iteratively restructuring workflow graphs by reordering nodes, modifying dependencies, and exploring alternative execution strategies guided by the task performance signal and convergence criteria.
@@ 5 Conclusion
We present EvoAgentX, an open-source platform that automates the generation, execution, evaluation and optimization of multi-agent workflows. It addresses key limitations of existing frameworks by eliminating the need for manual workflow design and providing support for dynamic, task-specific optimization. By integrating multiple optimization algorithms, including TextGrad, AFlow, and MIPRO, EvoAgentX enables the iterative refinement of agent prompts, tool configurations, and workflow topologies with minimal manual intervention. Experiments on diverse benchmarks, such as HotPotQA, MBPP, MATH and GAIA, demonstrate that EvoAgentX consistently achieves substantial performance improvements in multi-hop reasoning, code generation, mathematical problem solving, and real-world multi-agent applications. In the future, we will extend EvoAgentX with more optimization algorithms, richer tool integration, and long-term memory to further enhance agent adaptability and contextual awareness. We also plan to explore advanced evolution strategies Fang et al. (2025), including MASS Zhou et al. (2025), AlphaEvolve Novikov et al. (2025), and Darwin Gödel Machine Zhang et al. (2025c), to advance the strategies of multi-agent optimization.